OpenStudy (anonymous):

A line passes through the point (0, –4) and has a slope of 3/2 What is the equation of the line?

OpenStudy (anonymous):

y = 3/2x - 4

OpenStudy (anonymous):

omg thanks

OpenStudy (anonymous):

Lol Yep, remember the slope-intercept form it gave the slope and the y-intercept so fill in the m and b in this equation; y = (m)x + (b)

OpenStudy (anonymous):

yeah thanks and that isn't one of the choices though

OpenStudy (anonymous):

Really? What are the options?

OpenStudy (anonymous):

A. 3x - 2y = -8 B. 6x + 4y = -8 C. 2x + 4y = -1 D. 3x - 2y = 8

OpenStudy (anonymous):

the last one, I believe >.<, if you were to turn the equation into the slope-intercept, it'd be that one >.<

OpenStudy (acxbox22):

oh it has to be in standard form y=3/2x-4 -3/2x+y=-4 multiply by -2 on both sides 3x-2y=8 that is your answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

acxbox is stalking me -_-

OpenStudy (anonymous):

haha

OpenStudy (acxbox22):

\(\Huge\color{green}{No~problem!}\)

OpenStudy (anonymous):

i'm going to post more haha

OpenStudy (anonymous):

kaykay

OpenStudy (acxbox22):

@jadeybells no why would i do that? i am providing @karendiaz123 a better explanation of the answer than you are

OpenStudy (acxbox22):

it just happens to coincide that you are on this question at the same time as me

OpenStudy (anonymous):

:/ yeah, okay. In two questions already.

OpenStudy (acxbox22):

yeah, because i want to make sure the user understands the answer and how i got it you are just throwing the answer out there without an explanation